# Partial Resilient Leader-Follower Consensus in Time-Varying Graphs

### v1

@@ IV-B Convergent Set Analysis
Let all conditions in Theorem 1 hold. Define $N_{1}$ and $N_{2}$ as the numbers of convergent followers when all adversaries $a\in\mathcal{A}$ share $q_{a}^{u}[t,k]=0$ and $q_{a}^{u}[t,k]=1$, respectively, with their out-neighbors $u\in\mathcal{N}_{a}^{o}[t]$, for all $k\in[0,f-1]$ and $t\in\mathbb{Z}_{\geq 0}$. Then $N_{1}\leq|\mathcal{F}_{\mathcal{C}}|\leq N_{2}$.
@@
Let $\mathcal{F}_{\mathcal{R}}[t]=\{i\in\mathcal{F}_{\mathcal{M}}\mid q_{i}[t]=1\}$. Note that $i\in\mathcal{M}$ strictly follows (1) to update its activation state $q_{i}[t,k]$ for all iterations. By 1, we know $q_{a}^{u}[t,k]\in\{0,1\}$ for any $a\in\mathcal{A}$. Then, for a given $\mathbb{G}$, $|\mathcal{F}_{\mathcal{R}}[t]|$ is determined by whether $a\in\mathcal{A}$ shares $q_{a}^{u}[t,k]=1$ or $q_{a}^{u}[t,k]=0$ in each $k\in[0,f-1]$. By the protocol (1), for any $i\in\mathcal{F}_{\mathcal{M}}$, $q_{i}[t]=1$ if and only if $\sum_{j\in\mathcal{N}_{i}[t]}q_{j}^{i}[t,f-1]\geq 2F+1$. Therefore, $|\mathcal{F}_{\mathcal{R}}[t]|$ at time $t$ is minimized when all adversaries share $q_{a}^{u}[t,k]=0$ for every $k\in[0,f-1]$ and maximized when they share $q_{a}^{u}[t,k]=1$ for every $k$. Hence, all $a\in\mathcal{A}$ sharing $q_{a}^{u}[t,k]=0$ and $q_{a}^{u}[t,k]=1$ $\forall k\in[0,f-1]$ and $\forall t\in\mathbb{Z}_{\geq 0}$ will minimize and maximize $|\mathcal{F}_{\mathcal{C}}|$, respectively. ∎
@@
We know that $\mathcal{G}^{1}[t]$ is strongly $(2F+1)$-robust with respect to $\mathcal{L}_{\mathcal{M}}$ and $\mathcal{V}^{1}[t]\subseteq\mathcal{M}$. Hence, by Lemma 1, every $i\in\mathcal{V}^{1}[t]\cap\mathcal{F}_{\mathcal{M}}$ satisfies $\sum_{j\in\mathcal{N}_{i}[t]\cap\mathcal{M}}q_{j}^{i}[t,f-1]\geq 2F+1$. Therefore, if all adversaries share $q_{a}^{u}[t,k]=0$ for all $k\in[0,f-1]$, only the followers in $\mathcal{V}^{1}[t]\cap\mathcal{F}_{\mathcal{M}}$ are guaranteed to have $q_{i}[t]=1$ at time $t$. It follows that any $i$ with $|\mathcal{T}_{i}^{1}|=\infty$ has $\lim\sup_{t\to\infty}q_{i}[t]=1$, and thus convergent by Theorem 1. By Lemma 3, this defines the smallest possible convergent set, yielding $\mathcal{F}_{\mathcal{C}}^{1}\subseteq\mathcal{F}_{\mathcal{C}}$.
@@ V Simulation Results
We evaluate the effectiveness of our method through two sets of simulations. First, we illustrate Theorem 1 by comparing the performance of the BP-MSR algorithm against those of other existing resilient consensus algorithms. Second, we empirically validate the set bounds in 1. For all simulations, we use the graphs in Figure 1, each with leaders $\mathcal{L}=\mathcal{L}_{\mathcal{M}}=\{1,2,3\}$ and a $1$-local $\mathcal{A}=\{0\}$. The adversary shares $x_{0}^{j}[t]=1000\cdot\sin\big((t+j)/5\big)$ with the normal followers $j\in\{4,5,6,7,8,9\}$. The initial states of followers are generated randomly in the interval [1000,1000][-1000,1000]. Leaders update their states using $f_{r}$, drawing the same random value in [1000,1000][-1000,1000] every $50$ time step. Code is available here. [^1] https://github.com/joonlee16/partial-leader-follower-consensus

### v2

@@ IV-B Convergent Set Analysis
Let all conditions in Theorem 1 hold. Define $N_{1}$ and $N_{2}$ as the numbers of convergent followers when all adversaries $a\in\mathcal{A}$ share $q_{a}^{u}[t,k]=0$ and $q_{a}^{u}[t,k]=1$, respectively, with their out-neighbors $u\in\mathcal{N}_{a}^{o}[t]$, for all $k\in\{0,\dots,f-1\}$ and $t\in\mathbb{Z}_{\geq 0}$. Then $N_{1}\leq|\mathcal{F}_{\mathcal{C}}|\leq N_{2}$.
@@
Let $\mathcal{F}_{\mathcal{R}}[t]=\{i\in\mathcal{F}_{\mathcal{M}}\mid q_{i}[t]=1\}$. Note that $i\in\mathcal{M}$ strictly follows (1) to update its activation state $q_{i}[t,k]$ for all iterations. By 1, we know $q_{a}^{u}[t,k]\in\{0,1\}$ for any $a\in\mathcal{A}$. Then, for a given $\mathbb{G}$, $|\mathcal{F}_{\mathcal{R}}[t]|$ is determined by whether $a\in\mathcal{A}$ shares $q_{a}^{u}[t,k]=1$ or $q_{a}^{u}[t,k]=0$ in each $k\in\{0,\dots,f-1\}$. By the protocol (1), for any $i\in\mathcal{F}_{\mathcal{M}}$, $q_{i}[t]=1$ if there exists $\bar{k}\in\{0,\dots,f-1\}$ such that $\sum_{j\in\mathcal{N}_{i}[t]\setminus\mathcal{A}}q_{j}^{i}[t,\bar{k}]+\sum_{a\in\mathcal{A}}q_{a}^{i}[t,\bar{k}]\geq 2F+1$. Therefore, $|\mathcal{F}_{\mathcal{R}}[t]|$ at time $t$ is minimized when all adversaries share $q_{a}^{u}[t,k]=0$ for every $k\in\{0,\dots,f-1\}$ and maximized when they share $q_{a}^{u}[t,k]=1$ for every $k$. Hence, all $a\in\mathcal{A}$ sharing $q_{a}^{u}[t,k]=0$ and $q_{a}^{u}[t,k]=1$ $\forall k\in\{0,\dots,f-1\}$ and $\forall t\in\mathbb{Z}_{\geq 0}$ minimize and maximize $|\mathcal{F}_{\mathcal{C}}|$, respectively. ∎
@@
We know that $\mathcal{G}^{1}[t]$ is strongly $(2F+1)$-robust with respect to $\mathcal{L}_{\mathcal{M}}$ and $\mathcal{V}^{1}[t]\subseteq\mathcal{M}$. Hence, by Lemma 1, every $i\in\mathcal{V}^{1}[t]\cap\mathcal{F}_{\mathcal{M}}$ has $\bar{k}\in\{0,\dots,f-1\}$ such that $\sum_{j\in\mathcal{N}_{i}[t]\cap\mathcal{M}}q_{j}^{i}[t,\bar{k}]\geq 2F+1$. Thus, if all adversaries share $q_{a}^{u}[t,k]=0$ for all $k\in\{0,\dots,f-1\}$, only the followers in $\mathcal{V}^{1}[t]\cap\mathcal{F}_{\mathcal{M}}$ are guaranteed to have $q_{i}[t]=1$ at time $t$. It follows that any $i$ with $|\mathcal{T}_{i}^{1}|=\infty$ has $\lim\sup_{t\to\infty}q_{i}[t]=1$, and thus convergent by Theorem 1. By Lemma 3, this defines the smallest possible convergent set, yielding $\mathcal{F}_{\mathcal{C}}^{1}\subseteq\mathcal{F}_{\mathcal{C}}$.
@@ V Simulation Results
We evaluate the effectiveness of our method through two sets of simulations. First, we illustrate Theorem 1 by comparing the performance of the BP-MSR algorithm against those of other existing resilient consensus algorithms. Second, we empirically validate the set bounds in 1. For all simulations, we use the graphs in Figure 1, each with leaders $\mathcal{L}=\mathcal{L}_{\mathcal{M}}=\{1,2,3\}$ and a $1$-local $\mathcal{A}=\{0\}$. The adversary shares $x_{0}^{j}[t]=1000\cdot\sin\big((t+j)/5\big)$ with the normal followers $j\in\{4,5,6,7,8,9\}$. The initial states of followers are generated randomly in the interval [1000,1000][-1000,1000]. Leaders update their states using $f_{r}$, drawing the same random value in [1000,1000][-1000,1000] every $50$ time steps. Code is available here. [^1] https://github.com/joonlee16/partial-leader-follower-consensus